### v1

@@ A.8.2 Mean average precision metric
The classical metric used for object detection tasks is the mean average precision (mAP). While mAP can refer to slightly different quantities depending on the challenge, they all usually rely on the computation of the area under a recall-precision curve. The mAP is analog to the receiver operating characteristic curve (ROC) but for object detection tasks. In practice, mAP computation works as follows. All object detectors are expected to produce a score for their detection and sort them for the full dataset. Each prediction is then flagged as a true detection or a false positive using a match criterion that is usually based on a threshold limit using the classical IoU (Sect. 2.1 and Eq. 9). From this ordered list of true and false detections, it is possible to construct a running precision and recall quantity. The constructed raw recall-precision curve can present local drops due to some high-confidence objects sometimes being flagged as false detections. Therefore, the curve is modified to respect $p(r)=\max_{\tilde{r}>r}p(\tilde{r})$. The average precision (AP) is then obtained by computing the integral over this interpolated precision-recall curve.
@@
Usually, the AP is computed for each class individually. All the AP values are then averaged to construct a more general mAP score. For some challenges, mAP is computed for a single IoU value and expressed as mAP@IoU. For others, multiple IoU thresholds can be used independently or together to construct an averaged mAP over a given IoU range. Overall, this metric has the advantage of considering the full sensitivity of a given detector, which better represents the full discrimination capability between the background and the objects to detect (Sect. 5.1).
A.8.3 Imagenet classification pre-training
Following the description in Redmon & Farhadi (2016), we build a network backbone as similar as possible to the darknet-19 architecture, which comprises 19 convolutional layers and a few pooling layers. The main structural difference is that we use GN instead of BN. We first train this backbone on the ImageNet-2012 dataset (ILSVRC-2012, Russakovsky et al. 2015) composed of roughly 1.2 million single-label images with 1000 possible classes. This dataset’s size and diversity make it useful to pre-train networks to identify generic low-level features that are likely relevant for various applications. Due to the missing BN regularization, we must compensate with a more diverse data augmentation, still following common augmentation recipes. We first trained our network at the typical $224{\times}224$ input resolution up to a loss plateau that achieved a Top-1 accuracy of 70.1% and a Top-5 accuracy of 89.4\$ on the validation dataset. We then further train the network at a $448{\times}448$ input resolution, which is closer to the desired input resolution for our object detectors that will use this model as pretraining. With access to finer features, the network achieved a Top-1 accuracy of 73.9 and a Top-5 accuracy of 91.9. For comparison, the classical darknet-19 architecture reaches a top-1 accuracy of 76.5% and a top-5 accuracy of 93.3%. The remaining difference is likely due to the absence of BN in our architecture, which results in a lack of appropriate regularization. When diverging slightly more from the original darknet-19 to include more regularization in another form, we can close the accuracy gap, for example, by adding multiple dense layers with a high dropout rate between the global average pooling and the output layer. The corresponding scripts with all the hyper-parameters and the testing notebook are available in the CIANNA git repository, along with example applications to images from the test dataset.
A.8.4 PASCAL-VOC object detection
@@
The PASCAL-VOC (PASCAL Visual Object Classes, Everingham et al. 2010) dataset was a yearly computer vision challenge that ran from 2005 to 2012 and that included an object detection task. This dataset is still commonly used as a benchmark to compare the performances of different object detectors. One classical approach with this dataset is to combine the training and validation dataset from the 2007 and 2012 editions onto a single training sample and to evaluate the performances on the 2007 public test dataset of 5000 images or the 2012 hidden test dataset using the submission server. With this construction, the training set comprises roughly 16500 images with target bounding boxes of 20 possible classes. Using our high-resolution pre-trained backbone on ImageNet, we can drop the last pooling and output layer and add three new convolutional layers with $3{\times}3$ filters plus GN. Finally, we add our detection layer. The input resolution is also changed for a $416{\times}416$ input, which results in a $13{\times}13$ output grid. We note that in the original YOLO-V2 architecture, there is an additional skip connection between the second last layer and the layer just before the last pooling so that the network can access higher-resolution features for its prediction. As stated in Appendix A.8.1, the absence of this skip-connect will likely affect detection performances. While we could completely emulate the default YOLO association process, we use some of our method refinements to overcome the few structural limits we face. For example, we use the ${\rm DIoU}$ as matching score, $S_{ar}=3$, $L^{{\rm fIoU}}_{low}=-0.1$, $L^{{\rm fIoU}}_{C_{g}ood}=-0.1$, $\alpha_{best}=0.05$, and difficult object flagging. With these adjustments and a more diverse data augmentation, we achieve an ${\rm mAP@0.5}$ of 75.29% on the PASCAL 2007 test dataset. At the same resolution, the original YOLO-V2 model achieves 76.8%. Considering the structural limitations of our network that prevent us from perfectly emulating the original YOLO-V2 architecture, this result is a strong demonstration of the efficiency of our YOLO-CIANNA method. The corresponding scripts with all the hyper-parameters and the testing notebook are available in the CIANNA git repository, along with example applications to images from the test dataset.
A.8.5 COCO object detection
@@
YOLO-CIANNA 22.2 40.1 22.2 4.1 22.9 38.3 20.8 30.3 31.3 4.4 34.5 55.4
The COCO (Common Objects in Context, Lin et al. 2014) dataset was also a yearly computer vision challenge that ran from 2014 to 2017, and that included an object detection task. Like PASCAL, the COCO is still commonly used as a benchmark to compare the performances of different object detectors. In the 2017 version, the training set comprises roughly 118000 images, the validation set comprises 5000 images, and the hidden test set comprises roughly 40000 images (accessible through a test server). For the detection task, the objects are labeled with over 1000 classes. We use the same network backbone with pretraining on ImageNEt, with the same set of hyper-parameters we used for PASCAL and a $416{\times}416$ input resolution. Our result on the test dataset obtained using the submission server is presented in Table 6 using all the COCO metrics and is compared with the result of the classical YOLO-V2. We first observe that our method is competitive with the classical YOLO-V2 detector on this dataset despite our structural limitations. In addition, these metrics indicate that our method retrieves more large objects and better reconstructs the objects it manages to detect. However, the lower ${\rm mAP@0.5}$ score indicates that we recover fewer objects overall. This is likely due to a poorer detection of small objects, confirmed by the small ${\rm AR_{S}}$ corresponding to recall at small scales. This result was expected, considering that we lack the ending skip connection to allow our method to preserve more resolved features at the detection layer. The corresponding scripts with all the hyper-parameters and the testing notebook are available in the CIANNA git repository, along with example applications to images from the test dataset.
Regarding computing performances, we can achieve 690 images per second at this resolution using an RTX 4090 GPU with mixed precision enabled. On a much lighter A2000 mobile GPU, we still reach above 60 images per second, meaning that our implementation is perfectly suited for real-time detection on modern entry-level hardware. We note that this level of performance is similar to the one achieved by the original YOLO-V2 darknet implementation if we consider the raw TFlops scaling between our GPU and the one they used in the corresponding paper.
Finally, we highlight that our result is far from the current top scores on the COCO dataset that use much stronger architectures, but it is not our scope to compete with them. The presented benchmarks validate the method for a given relatively light architecture. Most of the time, the architectures that are truly useful for astronomical datasets are quite shallow due to much less feature diversity and complexity, the main challenges being placed in other aspects of the method design.
@@ Appendix B Classical YOLO-V2 backbone for SDC1
When first exploring the suitability of a given machine learning method to a new problem, the classical approach is to look at how this method was applied to a similar problem. This is especially true regarding network architectures and hyperparameters since exploring these aspects from scratch without prior knowledge of what type of architecture might be efficient can be extremely costly in terms of time and computing resources. In addition, if the new application is sufficiently close to other well-known applications of the same method, you can use pre-trained models to save some training time or improve the performance of your application. The global idea is that time and computing resources have already been invested in searching for the optimum architecture and hyper-parameters. We note that it is the approach used in Sortino et al. (2023) when testing the capability of a YOLO architecture on a source detection problem and that they concluded on improved results when using a pre-trained model. In fact, achieving apparently good results with known architectures or with pre-trained models does not guarantee that it is suited for the new application. In worst cases, it can even lead to the false conclusion that a specific approach usually tied to a given architecture is inefficient or unsuited for a specific application. To participate in this global discussion, we explore in this section the results we obtained on the SDC1 dataset using the classical YOLO-V2 darknet-19 architecture in a naive way and compare it to our reference result with our custom architecture.
@@ B.1 Selection of architectures and training setups
While the classical YOLO-V2 formalism could be emulated with our implementation, doing a thorough and documented comparison of the individual effects of all the modifications we made one by one would be too complicated. Here, we preserved all the data preparation, training setup, hyper-parameters selection, detection units setup, and prediction pipeline construction that we defined in Sect. 3 and only modified the network backbone and the starting weights. As a consequence, the setup of the last layer remains identical to our reference setup. To achieve the desired 16 reduction factor, we needed to cut a part of the darknet-19 architecture due to its reduction factor of 32. The classical way to do that is to preserve the first layers before the reduction factor reaches 32. Applying this approach to the architecture we used to emulate darknet-19 for the classical VOC datasets in the previous appendix section allows us to preserve the first 13 convolutional layers.
To provide interesting comparative results, we used two slightly different architectures. The first one is similar to using the pre-trained ImageNet network and modifying it for object detection. It is done by adding three new sets of convolutional layers with 1024 filters of size $3{\times}3$ followed by a GN, starting after our cut in the darknet-19 architecture, and then adding our detection layer. The only differences are the location of the pre-train architecture cut and the fact that we use the COCO model as our pre-trained network, which was already a detection task. We refer to this first architecture as YOLO-V2-base. The second one uses the same pre-trained part but uses the head of our adapted architecture consisting of the last 4 layers in Fig. 11 instead of the typical YOLO head. This structure has a more progressive increase in size and includes dropout regularization. We refer to this second architecture as YOLO-V2-regul. It is important to note the YOLO-V2-regul architecture is composed of 14.4 million parameters, which is very similar to our custom architecture. In contrast, the YOLO-V2-base is almost twice the size with a total of 28.3 million parameters, which is explained by the differences in the last 4 layers. In terms of computing performances, the YOLO-V2-regul architecture has almost the same performances as our custom architecture, while the YOLO-V2-base is around 20% slower.
We performed two independent training for each architecture, the first using the pre-trained weights for the corresponding layers and the second from scratch with classical random weight initialization for all layers. It means that we have a total of four independent models to compare. We use the same prediction pipeline with the parameters described in Sect. 3.9 still with a $512{\times}512$ input resolution for the four of them.
@@ B.2 Results and discussion
Table 7: . List of SDC1 scores and related properties for source catalogs from different architectures.
Method $M_{s}$ (Score) $N_{det}$ $N_{match}$ $N_{false}$ $N_{bad}\in N_{false}$ Purity $\bar{s}$
YOLO-V2-regul-scratch 445093 688112 645626 42486 15070 93.83% 0.7552
YOLO-V2-base-pretrain 423077 665304 620685 44619 16003 93.29% 0.7535
YOLO-V2-regul-pretrain 411541 642153 600193 41960 13398 93.47% 0.7556
YOLO-V2-base-scratch 386395 644371 591039 53332 15020 91.72% 0.7440
@@
The best score results for each of the four models are presented in Table 7. The best of these four models has a score $7\%$ lower than the best score achieved by our custom architecture, the latter reaching better scores after only 400 iterations (Fig. 13). We stress that the difference between these scores is always significantly larger than the score variability we observed over multiple training of the same setup with our reference architecture. This indicates that the observed ordering between these models is significant. We observe that for the YOLO-V2-base architecture, the pretraining improves the detection results, while it degrades it for the lighter YOLO-V2-regul architecture. The apparent flux distribution for these four models is very similar to our reference model. The main difference between these models is their completeness at a given purity, allowing some of them to detect more sources. We also note that these models can be ordered identically by their flux characterization score, which is coherent as the predicted flux accuracy is one of the match criteria.
@@
Our interpretation of these results is that the truncated darknet-19 architecture is not adapted to our application and encapsulates too much expressivity to be properly constrained by our training data. It appears that the pretraining does not encapsulate relevant information but rather acts as a startup regularization by reducing the accessible parameter space, preventing the network from over-fitting. The training of the YOLO-V2-base-scratch model is very unstable as we observe variations of about 100000 points between successive control steps when continuing the training after the best score iteration, which is a striking demonstration of overfitting. This model exhibits the poor results of an unconstrained network, while the YOLO-V2-regul-scratch demonstrates that adding a small amount of structural regularisation solves this issue better than pretraining. In our opinion, the strongest demonstration of the irrelevance of the features learned on COCO for our problem is that the YOLO-V2-regul-scratch model significantly outperforms the YOLO-V2-regul-pretrain model. In addition, we tested a variety of alternative setups around our YOLO-V2-regul architecture by changing the number of layers that start from their pre-trained weights. From this, we observe that the best achievable score degrades as the number of pre-trained layers increases. In this regard, we consider the common empirical observation about obtaining better results using pre-trained models as incomplete and possibly misleading regarding the relevance of low-level features learned from everyday-life images for astronomical datasets (e.g., Sortino et al. 2023; Wu et al. 2019). In addition, pre-trained models do not necessarily train faster than the others on this specific application, which is another sign of the non-relevant features they are starting from.
@@
Finally, achieving a better result with our custom backbone indicates that architectures that have been proved efficient on everyday-life images are not necessarily arranged in a way that is suited for extracting information in astronomical images. A striking difference between the two architectures is the absence of inter-layer normalization. We already identified in Sect. 3.6 that using GN between almost all layers in our custom architecture significantly reduces our best achievable score because it alters the absolute value of the input pixels, which contains information about the flux and the dynamic. Still, the presence of GN is not enough to explain the lower result of the YOLO-V2-regul-scratch model, highlighting that other architectural design choices matter.
@@ Appendix C Alternative training area for the SDC1
Figure 27: 2D histogram of the central coordinates of the sources from the full True catalog that passes our selection function. The red box indicates our modified SDC1 training area.
Figure 28: Source flux distribution histogram in log scale and using log-bins for the True catalog of the training region and the training catalog from our modified training area after the selection function.
Table 8: . List of scores and related properties based on the alternative training area for source catalogs from different architectures.
Method $M_{s}$ (Score) $N_{det}$ $N_{match}$ $N_{false}$ $N_{bad}\in N_{false}$ Purity $\bar{s}$
@@
YOLO-CIANNA-alt-train-ref 484145 708954 672906 36048 14988 94.91% 0.7731
Figure 29: Distribution of various kinds of ”objects” over the full SDC1 field, minus our alternative training area when necessary, for the YOLO-CIANNA-alt-train-ref model. The match and false detections are based on the scorer, while the missed ones are based on our selection function. The central coordinates of the field are RA=0 deg, Dec=-30 deg.
We exposed in Sect. 3.5 the issues that arise from having access to only a small sub-part of the center of the full image to train our detector. Due to the convolution with the primary beam, the flux sensitivity varies in the image field, and this variation is impossible to represent with only the provided training region. We mitigated this issue by adding two regions of the image that we consider empty in our training sample to provide examples of noise-only fields. The results presented in Sect. 4 demonstrated that this approach allows the detector to perform relatively homogeneously on the whole image. This section explores how the result varies when using a training area more suited for supervised learning approaches regarding the image field variation.

### v2

@@ A.8.2 Mean average precision metric
The classical metric for object detection tasks is the mean average precision (mAP). While mAP can refer to slightly different quantities depending on the challenge, they all usually rely on the computation of the area under a sensitivity curve, representing the precision as a function of recall. The mAP is analog to the receiver operating characteristic curve (ROC) but for object detection tasks. In practice, mAP computation works as follows. All object detectors are expected to produce a score for their detection and sort them for the full dataset. Each prediction is flagged as a true or false detection using a match criterion based on a classical IoU threshold limit (Sect. 2.1 and Eq. 9). From this ordered list of true and false detections, it is possible to construct a running precision and recall quantity. The constructed raw recall-precision curve can present local drops due to some high-confidence objects being flagged as false detections. Therefore, the curve is modified to respect $p(r)=\max_{\tilde{r}>r}p(\tilde{r})$. The average precision (AP) is then obtained by computing the integral over this interpolated precision-recall curve.
@@
Usually, the AP is computed for each class individually. All the AP values are then averaged to construct a more general mAP score. For some challenges, mAP is computed for a single IoU value and expressed as mAP@IoU. For others, multiple IoU thresholds can be used independently or together to construct an averaged mAP over a given IoU range. Overall, this metric has the advantage of sampling the full sensitivity of a given detector, corresponding to its discrimination capability between the background and the objects to detect (Sect. 5.1).
A.8.3 PASCAL-VOC object detection
@@
The PASCAL-VOC (PASCAL Visual Object Classes, Everingham et al. 2010) dataset was a yearly computer vision challenge that ran from 2005 to 2012, and that included an object detection task. This dataset is still commonly used as a benchmark to compare the performances of different object detectors. One classical approach is to combine the training and validation dataset from the 2007 and 2012 editions onto a single training sample comprising around 16500 images containing multiple bounding boxes belonging to 20 possible classes. The evaluation of the detector performances is then done on the 2007 test dataset of 5000 images. Using our high-resolution pretrained backbone on ImageNet, we can drop the last pooling and output layer and add three new convolutional layers with $3{\times}3$ filters plus group normalization. Finally, we add our detection layer. The input resolution is changed for a $416{\times}416$ input, which results in a $13{\times}13$ output grid. Regarding the setup of our advanced association function, we use the ${\rm DIoU}$ as our matching score, $S_{ar}=3$, $L^{{\rm fIoU}}_{\rm low}=-0.1$, $L^{{\rm fIoU}}_{C}=-0.1$, $\alpha_{\rm best}=0.05$, and some difficult object flagging. After training this network for a few hundred epochs, we achieve an ${\rm mAP@0.5}$ of 74.98 on the PASCAL 2007 test dataset. At an identical resolution, the original YOLO-V2 model achieves 76.8. Considering the structural limitations of our approximated darknet-19 backbone, this result demonstrates the efficiency of our YOLO-CIANNA method.
A.8.4 COCO object detection
@@
YOLO-CIANNA 21.9 40.1 21.8 4.1 22.8 37.9 20.7 30.2 31.2 4.5 34.6 55.0
The COCO (Common Objects in Context, Lin et al. 2014) dataset was also a yearly computer vision challenge that ran from 2014 to 2017, and that included an object detection task. Like PASCAL, COCO is still commonly used as a benchmark to compare the performances of different object detectors. In the 2017 version, the training set comprises roughly 118000 images, the validation set comprises 5000 images, and the hidden test set comprises roughly 40000 images (accessible through a test server). For the detection task, the objects are labeled with over 1000 classes. We use the same network backbone pretrained on ImageNet, with the same set of hyperparameters we used for PASCAL and a $416{\times}416$ input resolution. Our result on the test dataset obtained using the submission server is presented in Table 6 with all the COCO metrics and is compared with the result of the classical YOLO-V2 model. We first observe that our method is competitive with the classical YOLO-V2 detector on this dataset despite our structural limitations. In addition, these metrics indicate that our method retrieves more large objects and better reconstructs the objects it manages to detect. However, the lower ${\rm mAP@0.5}$ score indicates that we recover fewer objects overall. This is likely due to a poorer detection of small objects, confirmed by the small ${\rm AR_{S}}$ corresponding to recall at small scales. This result was expected, considering that we lack the ending skip connection to allow our method to preserve more resolved features at the detection layer.
Regarding computing performances, we can achieve 690 images per second at this resolution using an RTX 4090 GPU with mixed precision enabled. On a much lighter A2000 mobile GPU, we still reach above 60 images per second, meaning that our implementation allows real-time detection on modern entry-level hardware. We note that this level of performance is similar to the one achieved by the original YOLO-V2 darknet implementation.
Finally, we highlight that our result is far from the current top scores on the COCO dataset, which require much stronger architectures, but it is not our scope to compete with them. The presented benchmarks validate the method for a given relatively light architecture. Most of the time, the architectures that are truly useful for astronomical datasets are relatively shallow due to much less feature diversity and complexity, the main challenges being placed in other aspects of the method design.
@@ Appendix B Classical YOLO-V2 backbone for SDC1
Table 7: . SDC1 scores and related properties for source catalogs from different models and architectures.
Method $M_{s}$ (Score) $N_{\rm det}$ $N_{\textrm{match}}$ $N_{\rm false}$ $N_{\rm bad}\in N_{\rm false}$ Purity $\bar{s}$
darknet-19-regul-scratch 445093 688112 645626 42486 15070 93.83% 0.7552
darknet-19-base-pretrain 423077 665304 620685 44619 16003 93.29% 0.7535
darknet-19-regul-pretrain 411541 642153 600193 41960 13398 93.47% 0.7556
darknet-19-base-scratch 386395 644371 591039 53332 15020 91.72% 0.7440
When exploring the suitability of a given method to a new problem, the classical approach is to look at how this method was applied to a similar problem. This is especially true for network architectures and hyperparameters since exploring these aspects from scratch can be extremely costly in terms of time and computing resources. If the new application is sufficiently close to other well-known applications of the same method, one could even use pretrained models to save some training time or improve the performance of their new model. For example, it is the approach used in Sortino et al. (2023) where they test the capability of a YOLO architecture on a source detection problem. They conclude that the results were improved when using a pretrained model. However, achieving apparently good results with known architectures or with pretrained models does not guarantee that the new application will have close to optimal results. In cases where it achieves bad results, it can even lead to the false conclusion that the whole approach is inefficient while only the backbone architecture might be unsuited for the task. To participate in this global discussion, we explore in this section the results we obtained on the SDC1 dataset using the classical YOLO-V2 darknet-19 architecture in a naive way and compare it to our reference result with our custom architecture.
@@ B.1 Selection of architectures and training setups
While the classical YOLO-V2 formalism could be emulated with our implementation, doing a thorough and documented comparison of the individual effects of all the differences would be too complicated. In this section, we preserved all the data preparation, training setup, hyperparameters selection, detection units setup, and prediction pipeline construction that we defined in Sect. 3 and studied variations in the network backbone and starting weights. To achieve the same reduction factor of 16 as our reference SDC1 model with pretrained darknet-19 backbones, it is necessary to cut the end part of the darknet-19 architecture before it reaches a reduction factor of 32.
We compare two slightly different architectures. The first one is obtained by taking the first 13 convolution layers of our darknet architecture pretrained on the COCO dataset and adding three new convolutional layers with 1024 filters of size $3{\times}3$, each followed by a group normalization. We then add our detection layer. This first architecture is referred to as darknet-19-base. The second one uses the same pretrained part but then adds layers to recreate the detection head structure of our custom architecture consisting of the last four layers in Fig. 11. This architecture has a more progressive increase in size and includes dropout regularization. This second architecture is referred to as darknet-19-regul. It is essential to note the darknet-19-regul architecture is composed of 14.4 million parameters, which is very similar to our custom architecture. In contrast, the darknet-19-base is almost twice the size with a total of 28.3 million parameters, which is explained by the differences in the last four layers. Regarding computing performances, the darknet-19-regul architecture has almost the same performances as our custom architecture, while the darknet-19-base is around 20% slower.
We performed two independent training for each architecture, the first using the pretrained COCO weights for the corresponding layers (-pretrain) and the second from scratch using classical random weight initialization for all layers (-scratch). Therefore, we have a total of four trained models to compare. We use the same prediction pipeline as our reference model for all of them, following the description in Sect. 3.9.
@@ B.2 Results and discussion
The best score results for each of the four models are presented in Table 7. We can first observe that the best one has a score $7\%$ lower than our reference model from Table 5, which reaches better scores after only 400 iterations (Fig. 13). We stress that the results are reproducible and stable over multiple training with different initial weight values. In addition, the difference in score between models is significant as it is higher than the typical score fluctuation observed over multiple training of the same setup. From these results, we see that for the darknet-19-base architecture, pretraining improves the detection results, while it degrades it for the lighter darknet-19-regul architecture. The apparent flux distribution for these four models is very similar to our reference model. The main difference between these models is their completeness at a given purity, allowing some of them to detect more sources. We also note that models that achieve the best score have the highest characterization accuracy. This is expected, as the predicted flux accuracy is one of the match criteria.
@@
Our interpretation is that the truncated darknet-19 architecture is not adapted to our application and encapsulates too much expressivity to be properly constrained by our training data. It appears that the pretrained weights do not encapsulate relevant information but still act as a startup regularization by reducing the accessible parameter space, preventing the network from over-fitting. The training of the darknet-19-base-scratch model is very unstable as we observe variations of about 100000 points between successive control steps when continuing the training after the best score iteration, which is a striking proof of overfitting. This model exhibits the poor results of an unconstrained network, while the darknet-19-regul-scratch demonstrates that adding a small amount of structural regularization solves this issue better than pretraining. In our opinion, the most robust demonstration of the irrelevance of the features learned on COCO for our problem is that the darknet-19-regul-scratch model significantly outperforms the darknet-19-regul-pretrain model. In addition, we tested a variety of alternative setups around our darknet-19-regul architecture by changing the number of layers that start from their pretrained weights, the rest being initialized from scratch. We observed that the best achievable score degrades as the number of pretrained layers increases. In this regard, we consider the common empirical observation about obtaining better results using pretrained models as incomplete and possibly misleading regarding the relevance of low-level features learned from everyday-life images for astronomical datasets (e.g., Sortino et al. 2023; Wu et al. 2019). In addition, pretrained models do not necessarily train faster than the others on this specific application, which is another sign of the nonrelevance of the features they are starting from.
@@
Finally, achieving better results with our custom backbone indicates that architectures that have been proven efficient for everyday-life images are not necessarily suited for astronomical images. A striking difference between the two architectures is the absence of inter-layer normalization. We already identified in Sect. 3.6 that using group normalization between most layers in our custom architecture significantly reduces our best achievable score because it alters the absolute value of the input pixels, which contains information about the source flux and the input dynamic. Still, group normalization is not enough to explain the lower result of the darknet-19-regul-scratch model (Sect. B), highlighting that other architectural design choices matter.
@@ Appendix C Alternative training area for the SDC1
Table 8: . Scores and related properties based on the alternative training area for source catalogs from different models.
Method $M_{s}$ (Score) $N_{\rm det}$ $N_{\textrm{match}}$ $N_{\rm false}$ $N_{\rm bad}\in N_{\rm false}$ Purity $\bar{s}$
@@
YOLO-CIANNA-alt-train 484145 708954 672906 36048 14988 94.91% 0.7731
Figure 27: Two-dimensional histogram of the central coordinates of the sources from the full True catalog that pass our selection function. The red box indicates our alternative training area.